Calculate the normality of the following:
a. 26.5 grams nitric acid in 2L of solution
b. 150g of calcium carbonate in 500 ml solution
1
Expert's answer
2020-09-04T05:33:59-0400

Solution:

Normality (N) It is the number of gram equivalent of solute present in one litre of solution.

  • Normality = Number of gram equivalents / Volume of solution in litres
  • Number of gram equivalents = Weight of solute / Equivalent weight of solute
  • N = Weight of Solute (gram) / [Equivalent weight × Volume (L)]

Part (a):

nitric acid = HNO3

To find Equivalent weight of HNO3, use the formula Eq = MW / n

  • Eq = Equivalent weight
  • MW = Molecular weight in grams/mole
  • n = number of equivalents

For every mole of nitric acid, there is one H+ ions, n = 1.

Molecular weight (MW) of HNO3 is 63.01 g/mol.

Consequently,

Eq = (63.01 g/mol) / 1 = 63.01 g/mol

Thus:

Normality of HNO3 = (26.5 g) / [63.01 g/mol × 2 L] = 0.21 N

Normality of HNO3 is 0.21 N


Answer (a): Normality of HNO3 is 0.21 N.


Part (b):

calcium carbonate = CaCO3

To find Equivalent weight of CaCO3, use the formula Eq = MW / n

Molecular weight (MW) of CaCO3 is 100.087 g/mol.

Valency (Ca2+; CO32-) is the n-factor for salt, n = 2.

Consequently,

Eq = (100.087 g/mol) / 2 = 50.0435 g/mol

Thus:

Normality of CaCO3 = (150 g) / [50.0435 g/mol × 0.5 L] = 5.995 N = 6N

Normality of CaCO3 is 6 N


Answer (b): Normality of CaCO3 is 6 N.